In a series RC circuit with an AC source, R = 300 Ω, C = 25 μF, ℰ^0 = 50 V and v = 50/π Hz. Find the peak current and the average power dissipated in the circuit.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. R = 300Ω, C = 25 μF = 25 * 10^-6 F, ε base 0 = 50 V, f = 50 Hz X base c = 1/ωc = 1/50/π *2π *25 *10^-6 = 10^4/25 Z = √R^2 + X base c^2 = √(300)^2 + (10^4/25)^2 = √(300)^2 + (400)^2 = 500 (a) Peak current = E base 0/Z = 50/500 = 0.1 A (b) Average Power dissipitated, = E base rms I base rms Cos ∅ = E base 0/√2 * E base 0/√2Z * R/Z = E base 0^2/2Z^2 = 50 * 50*300/2* 500 *500 = 3/2 = 1.5 ω.